Calculate the columb force (F1), between two equal charges placed in vacuum separated at a distance square of r meters apart. By what factor does F1 changes when the separation is doubled?
1
Expert's answer
2019-04-17T11:08:18-0400

The Coulomb force (F1) between the two equal charges separated by a distance r from each other in vacuum can be calculated as follows:


"F_1 = \frac{k q^2}{r^2},"

where k = 9×109 N·m2·C−2 is the Coulomb constant.

If the separation between the charges is doubled, i.e.


"R = 2r"

then the Coulomb force reduces as follows:


"F_2 = \frac{k q^2}{R^2} = \frac{k q^2}{4r^2}=\frac{F_1}{4}"

Answer: reduction by a factor 4.